# Testing the heavy decaying sterile neutrino hypothesis at the DUNE near detector

### v1

@@ ACKNOWLEDGMENTS
We thank Guadalupe Moreno-Granados for collaboration at an early stage of this project. The work of S.S.C. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project No. 510963981. S.S.C. also acknowledges financial support from the LabEx P2IO (ANR-10-LABX-0038 - Project “BSMNu”) in the framework of the “Investissements d’Avenir” (ANR-11-IDEX-0003-01) managed by the Agence Nationale de la Recherche (ANR), France. The work of S.L. is supported in part by the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No. 860881-HIDDeN. S.L. also acknowledges the hospitality of the Kavli IPMU while working on this project. O. G. M. was supported by the CONAHCyT Grant 23238 and by SNII-Mexico. O. G. M. would like to thank Luis Delgadillo for useful discussions.

### v2

@@ Acknowledgments
We thank Guadalupe Moreno-Granados for collaboration at an early stage of this project. The work of S.S.C. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project No. 510963981. S.S.C. also acknowledges financial support from the LabEx P2IO (ANR-10-LABX-0038 - Project “BSMNu”) in the framework of the “Investissements d’Avenir” (ANR-11-IDEX-0003-01) managed by the Agence Nationale de la Recherche (ANR), France. The work of S.L. is supported in part by the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie Staff Exchange grant agreement No. 101086085 – ASYMMETRY. S.L. also acknowledges the hospitality of the Kavli IPMU while working on this project. O. G. M. was supported by the CONAHCyT Grant 23238 and by SNII-Mexico. O. G. M. would like to thank Luis Delgadillo for useful discussions.